Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
hui, gouaty, gherardini - 17:42 Wednesday 14 February 2024 (63269) Print this report
OMC mode matching fine tuning

Starting from ITF in Low Noise 3.

B1_DARM_OG is hitting 1.5. while DCP is around 250 Hz.

Reference position for SDB1: LC_Z = 3789 um, F0_X = 2600 um

 

1/ Scan of the SDB1 bench longitudinal position in Low Noise 3:

We started by scanning the SDB1 bench longitudinal position by acting on the suspension top stage (F0_X).

Fig.1 shows the evolution of the figures of merit during the scan of the longitudinal position of SDB1. It seems that moving the bench towards F0_X = 3400 has improved the mode matching because the B1s signal demodulated at 491 Hz is going towards zero. However this interpretation may not be correct as SR alignment is also changing at the same time as we are translating the bench, and the optical gain is correlated with the fluctuations of the DCP frequency. It seems that LN3 is not a configuration where we can easily perform this tuning (Low Noise 2 may be better for this purpose).

For this reason, we decided to unlock the ITF. Before unlocking, we tried to move the meniscus lens by 100 steps, which unlocked the ITF immediately at 09h11m20 utc.

Since we cannot move the meniscus lens when the ITF is locked, we decided to perform the coarse tuning of the mode matching with the NI single bounce beam. We asked Fabio to misalign PR TY further to bring the ghost beam out of the B1p camera.

We removed the B5 QD2 offsets used for dark fringe operations at 09h19m17 utc.
 

2/ OMC scan to find the mode matching

We performed an OMC scan at 09h24 utc (Fig.2): We found 11.7 uW on B1_PD3 TEM00, order 2 < 0.25 uW, but misalignment (order 1) is significant (~2.5 uW).

We locked the OMC on the TEM00 at 09h36 utc and adjusted the alignment with B5 QD2 offsets. The best alignment of the OMC is obtained with the following offsets: B5_QD2_H=-130, B5_QD2_V=+100.

We performed another scan of the OMC starting at 10h11m20 utc (Fig.3): TEM00 13.5 uW, order 1: 0.1 uW (0.7 %), order 2: 0.75 uW (residual mode mismatch = 5.6%)
 

3/ Scan of SDB1 and meniscul lens position

We locked again the OMC at 10h22m20 utc, and then we started to translate the bench to improve further the mode matching. We used as observables the OMC transmitted power (B1_PD3_DC) and the transmitted power normalized by the reflected power.

OB_F0_X is first changed from 3400 to 4000 um. We paused the scan to realign the OMC with the picomotors on the folding mirrors: OMC_TY_1: -200 steps and OMC_TY_2: -400 steps ; OMC_TX_1 : 150 steps, OMC_TX_2: 300 steps

We continued the scan of OB_F0_X, reaching OB_F0_X =4600, which is close to the end of the range. We realigned the bench.

Starting moving the meniscus lens L1_Z from 11h06 utc:

L1_Z: -2500 -1000 -2500 -2500 -2500 -2500 > Total: -13500 steps

BENCH_TX: -2500, -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 > Total : -22500  steps

L1_Z (11h23): -2500 -2500 -2500 -2500 -2500 -2500 > Total : -15000 steps

BENCH_TX: -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 > Total : -20000 steps

L1 Z (11h32): -2500 -2500 -2500 -2500 -2500 -2500 > Total : -15000 steps

BENCH_TX: -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 > Total : -25000 steps

We realigned the bench and observed a clear improvement of the transmitted power normalized by B1s.

Continued moving L1_Z:

L1_Z (11h49) : -2500 -2500 -2500 -2500 -2500 -2500 > Total : -15000 steps

BENCH TX : -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500  -2500 -2500 > Total : -25000 steps

L1_Z (12h46):  -2500 -2500 -2500 -2500 -2500 -2500 > Total : -15000 steps

BENCH TX : -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 > Total : -25000 steps

L1 Z (12h51) : -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 > Total : -20000 steps

BENCH_TX : -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 > Total : -20000 steps

We then realigned the OMC with picomotors.

L1 Z (13h12) : -2500  -2500 -2500 -2500 -2500 -2500 -2500 -2500 > Total : - 20000 steps

BENCH_TX : -2500  -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 -2500 > Total : -25000 steps

We think that we have passed the optimum of transmission.

Total number of steps with L1_Z: -113500 steps.

Total number of steps with BENCH TX : -162500 steps.

Thus we brought back the bench in Z by steps of 500 um, up to the initial position: OB_F0_X = 2600

It seems that we are close to the optimum transmission after the bench position was restored.

The evolution of the OMC throughput during the mode matching tuning in single bounce is shown on Fig.4.

4/ New OMC scan after optimization of mode matching

Scan of the OMC at 14h22 utc (Fig.5): 13.9 uW on TEM00, order 1 = 0 (seems lower than 0.2%), order 2 = 0.33 uW > this corresponds to 2.4% of residual mode mismatch. The image of the second order mode is a nice four-leaf clover.

At the end of the shift we restored the B5 QD2 offsets used for dark fringe operations.


Conclusion : We have completed the coarse tuning of the OMC mode matching in single bounce by adjusting the distance between the SR mirror and the meniscus lens. The remaining step will be to check in Low Noise 2 if some fine tuning needs to be done to optimize further the optical gain.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×